OpenStudy (starryknight):

Determine the equation of the axis of symmetry. y = -5x2 - x + 9 x = 3 2 x = 1 -30 x = -1 10 x = 1 10

OpenStudy (anonymous):

the line of symmetry is found @ x=-b/2a x=-(-1)/2(5)=1/10.......ANS

OpenStudy (anonymous):

Sry i made a mistake the answer is x=-1/10 bcouse a=-5